# Modelling the evolution and influence of dust in cosmological simulations that include the cold phase of the interstellar medium

### v1

@@ 2.5 Grain size evolution
Other processes have been identified that conserve the mass in dust, but modify the grain size distribution. Despite not affecting the grain mass directly, this has a knock-on effect on the evolution of grains, with the accretion and sputtering timescales being sensitive to the size of grains.
@@ 2.5.1 Shattering
where the timescale normalisation value assumes a material density of 3 ${\rm g\;cm^{-3}}$ for grains, and $a_{L}$ is the radius of large grains (0.1 µm in our model). The Granato et al. (2021) prescription differs from previous applications in that the shattering does not shut off abruptly above a gas density of $n_{\rm H}=1\;{\rm cm^{-3}}$, but instead increases more gradually above this density. Note, as in sputtering, we are using the unmodulated density, $n_{\rm H}$ (as opposed to $n_{\rm H}^{\prime}$), as sputtering is a diffuse-gas process.
@@ 2.5.2 Coagulation
[…]
Table 2: Simulation variations used in this work. All runs represent a 253 Mpc3 cosmological box, initialised with 3763 gas particles of mass $m_{\rm g}=1.8\times 10^{6}\;{\rm M_{\odot}}$ (and 4 times as many dark matter particles, with mass $m_{\rm DM}=2.4\times 10^{6}\;{\rm M_{\odot}}$). We adopt ‘m6’ as shorthand for this resolution (from the COLIBRE parlance). Simulations are run using the COLIBRE galaxy formation model (Schaye et al. in prep.), using the swift code (Schaller et al., 2024, see section 3.1). The columns from left to right indicate: 1) the simulation name, 2) if the dust model is coupled to the cooling (see § 3.5), 3) if the model includes small grains rather than large, if we include processes of 4) accretion, 5) destruction 6) or size transfer, 7) the maximum $C$ value 8) if $C$ has a fixed value 9) the size representing the small grain bin 10) the chemical group assumed for silicates and 11) the diffusion constant (§ 2.6).
1) 2) 3) 4) 5) 6) 7) 8) 9) 10) 11)
(a) [µm]
Fiducial 100. 0.01 Olivine 0.01
FidUncoupled 100. 0.01 Olivine 0.01
SeedOnly n/a 0.01 Olivine 0.01
NoDestruction 100. 0.01 Olivine 0.01
LargeGrainsOnly 100. 0.01 Olivine 0.01
MaxC10 10. 0.01 Olivine 0.01
constC30 30. 0.01 Olivine 0.01
NoC 1. 0.01 Olivine 0.01
PyroSil 100. 0.01 Pyroxene 0.01
SmallerGrains 100. 0.005 Olivine 0.01
NoDiff 100. 0.01 Olivine 0
LoDiff 100. 0.01 Olivine 0.001
HiDiff 100. 0.01 Olivine 0.1
(a) i.e. the clumping factor
@@ 2.6 Grain transport & diffusion
3 Simulations & The COLIBRE model
The dust evolution model we described is integrated into the COLIBRE model of galaxy formation physics (Schaye et al in prep.), run using the gravity and hydrodynamics solver SWIFT (Schaller et al., 2024). In the multi-resolution parlance of COLIBRE, we use m6 to refer to a gas (DM) particle mass of $1.84\times 10^{6}\;{\rm M_{\odot}}$ ($2.42\times 10^{6}\;{\rm M_{\odot}}$). We note that COLIBRE employs 4$\times$ more dark matter particles than gas, achieving a relatively high DM resolution. In this section, we first provide technical details of the simulation runs and parameter variations used in this work for context (§3.1), and then describe the general COLIBRE model, including the coupling between dust physics and baryon physics more broadly.
3.1 Simulations
Having described our dust model above (§2), we describe the simulations used in this work and our choice of parameter variations. Throughout this work we assume the $3\times 2$ pt cosmology, combining all data and constraints presented with the DES Y3 results (Abbott et al., 2022); $h=0.681$, $\Omega_{\rm m}=0.306$, $\Omega_{\rm b}=0.0486$, $\sigma_{8}=0.807$, $n_{\rm s}=0.967$.
We use cosmological volume initial conditions for the simulations in this work. This choice allows us to evolve a population of galaxies in a cosmological context, and to inspect galaxy scaling relations. This is in contrast to the idealised disc simulations often used in the literature. While idealised galaxy simulations allow a more controlled comparison where the ‘same’ galaxy is run each time (with modified parameters), the set up of such runs is somewhat artificial; the disc component should co-evolve with the dust and metal content of galaxies, and building enrichment up in an already formed disc is unphysical. While such simulations can show how the dust and metals are maintained in galaxies, the build up of enrichment requires cosmological simulations.
We focus on a $25^{3}\;{\rm cMpc^{3}}$ volume, with an initial $376^{3}$ gas particles of $m_{\rm gas}=1.8\times 10^{6}\;{\rm M_{\odot}}$ and $4\times 376^{3}$ dark matter particles with $m_{\rm DM}=2.4\times 10^{6}\;{\rm M_{\odot}}$, representing matched (4$\times$ higher) baryonic (dark matter) resolution relative to fiducial EAGLE (Schaye et al., 2015).
We identify collapsed structures using the freinds-of-friends (FoF) algorithm post processed using the publically available HBT-HERONS subhalo finder described by Forouhar Moreno et al. (2025) to assign particles to self-bound substructures. This builds on the HBT+ code for $N$-body simulations (Han et al., 2018), with critical improvements for the hydrodynamical context, considering the histories and hierarchical formation of subhalos. The subhalo locations and, optionally, the particles are then taken up by the SOAP tool (McGibbon et al., 2025), computing a wide range of aggregated properties using a range of 3D and projected apertures of varying physical size.
The simulations run for analysis in this work and their defining parameters are listed in Table 2. The Fiducial run uses the full dust model described above, and also includes the effects of "coupling” our dust model to gas cooling, heating and phase transition processes, detailed in §3.6. Our runs incorporate dust coupling by default (excepting SeedOnly where dust fractions are too unphysical), with an uncoupled counterpart to isolate the influence of the dust evolution model on the cooling, e.g. FidUncoupled. The SeedOnly and NoDestruction simulations isolate evolutionary processes. Size transfer effects are tested by LargeGrainsOnly. A number of runs are dedicated to exploring the influence of parameters; particularly a subgrid clumping factor, $C$, (described in §3.7; NoC, ConstC30 and MaxC10) and the turbulent diffusion constant, $C_{\rm d}$ (see §2.6; NoDiff, LoDiff and HiDiff). We also present runs exploring pyroxene silicate chemistry, PyroSil, and the representative size of small grains, SmallerGrains. The use of coupled cooling physics by default means we see the impact of these dust model changes on the physics and resultant galaxy populations. We will explore the most pertinent variations alongside our primary low-redshift results in §4, as well as presenting some additional variations in Appendix B.
3.2 Energy feedback
[…]
For the thermal part, the target heating temperature, $\Delta T_{\rm SN}(n_{\rm H})$, is an increasing function of local gas density with resolution-dependent floor and ceiling values ($76.5<\log_{10}\Delta T_{\rm SN,min}/{\rm K}<$ 7 and $87.5<\log_{10}\Delta T_{\rm SN,max}/{\rm K}<$ 8) calibrated from test runs. This approach mitigates overly destructive feedback and extensive bubbles in low-density gas. Finally, we adopt the statistically isotropic gas particle particle selection for energy injection of Chaikin et al. (2022), with the thermal feedback also using a variable fraction of energy per single CCSN in units of $10^{51}~{}\mathrm{erg}$. This is subject to a variable coupling efficiency, $f_{\rm E}$, to avoid over-cooling at high density (Crain et al., 2015). In COLIBRE this is implemented as a sigmoid in SSP birth pressure, calibrated to reproduce the $z=0$ galaxy stellar mass function and galaxy sizes (Chaikin et al. in prep.).
Alongside stars, COLIBRE accounts for feedback by supermassive black holes (SMBHs). SMBH particles are first seeded for friends-of-friends halos (identified on-the-fly) in a resolution-dependent way, using an SMBH particle of fixed mass ($3\times 10^{4}\;{\rm M_{\odot}}$ at m6), for halos above a halo FoF mass ($10^{10}\;{\rm M_{\odot}}$ at m6). These particles can then accrete gas, merge and feed energy back to the surrounding media. The COLIBRE model includes two prescriptions for AGN feedback: purely thermal energy injections and a hybrid model combining thermal and kinetic jet feedback. In this work, we employ the purely thermal feedback prescription. Following similar arguments for effective feedback in CCSN, feedback is quantised to heat surrounding gas by a target temperature increment, $\Delta T_{\rm AGN}$, to mitigate numerical cooling losses. SMBHs build up an energy reservoir until enough is accrued for a single feedback event. The energy for feedback comes from gas accretion; mass energy is converted to a luminosity, that couples back to the gas with fixed efficiency ($5\%$ at m6 resolution). In COLIBRE, $\Delta T_{\rm AGN}$ is a function of black hole mass, again enforcing floor and ceiling values ($\log_{10}\Delta T_{\rm AGN,min}/{\rm K}=6.5$ and a resolution-dependent $109.5<\log_{10}\Delta T_{\rm SN,max}/{\rm K}<$ 10) helping to maintain a reasonable sampling and regularity of BH feedback across the mass range of AGN. Feedback is meted out to the gas particle closest to the SMBH at feedback time.
3.3 Star formation
3.4 Stellar mass loss and chemical enrichment
Enrichment timesteps for star particles are designed to balance accurate sampling of stellar evolution within each SSP against computational cost. To capture the early, most active phase of enrichment, we use fine timesteps ($\leq 1$ Myr) for SSPs younger than 40 Myr. For older populations ($>100$ Myr), the time intervals after which mass transfer is performed increase proportionally to age.
3.5 Coupling dust to COLIBRE physics
For P25 dust, it is assumed that grains are immediately destroyed at $T>10^{5}\;{\rm K}$ and do not exist in relatively diffuse gas, represented by a minimum hydrogen column density, ($N_{\rm min}=3.1\times 10^{15}\;{\rm cm}^{-2}$), such that $\mathcal{DTG}_{\rm hyb}(T>10^{5}\;{\rm K})=0$ and $\mathcal{DTG}_{\rm hyb}(N_{\rm H}\leq N_{\rm min})=0$. At high column densities ($N_{\rm H}>10^{20}\;{\rm cm}^{-2}$), a constant dust-to-metal ratio is assumed, such that [^6] As in P25, we use the $Z_{\odot}=0.0134$ value of Asplund et al. (2009) where appropriate. $\mathcal{DTG}_{\rm hyb}\propto Z_{\rm gas}/{\rm Z_{\odot}}$. For intermediate column densities ($3.1\times 10^{15}<N_{\rm H}/{\rm cm}^{-2}<10^{20}$), an additional taper is introduced to smooth the transition between the metallicity-dependent $\mathcal{DTG}_{\rm hyb}(N_{\rm H}=10^{20}\;{\rm cm}^{-2})\propto Z_{\rm gas}/{\rm Z_{\odot}}$ and $\mathcal{DTG}_{\rm hyb}(N_{\rm H}=3.1\times 10^{15})=0$, assuming a power law in column density with index $\alpha=1.4$. This is chosen based on the Kennicutt-Schmidt relation (Kennicutt, 1998), and assuming a scaling between the star formation rate density and dust production. At high density and solar metallicity, the $\mathcal{DTG}$ is normalised to a representative MW value of $\mathcal{DTG}_{\rm MW}=6.6\times 10^{-3}$.
where $\nu_{\rm HI}$ and $n_{\rm HI}$ are the thermal velocity and number density of atomic hydrogen, $\epsilon_{\rm H_{2}}$ is the recombination efficiency of molecular hydrogen, $S_{\rm H}$ is a dimensionless sticking coefficient and $n_{\rm d}\sigma_{\rm d}$ is the number density of dust grains times the grain cross section, i.e. the total cross section per unit volume. While $S_{\rm H}$ and $\sigma_{\rm d}$ are temperature-dependent, a constant dust temperature of 10 K is assumed [^7] The H2 formation rate was shown to be approximately constant for $6\;{\rm K}\lessapprox T_{\rm dust}\lessapprox 50\;{\rm K}$ in fig. A1 of Richings et al. (2014a).. It is then the $n_{\rm d}\sigma_{\rm d}$ term that depends on the local mass fraction and size distribution of grains, with the other terms depending on the material properties and local gas properties. Dust grains can also contribute to catalysis of ion recombination via available dust surfaces (see e.g. Weingartner & Draine, 2001), and transfer energy between the dust and gas (Richings et al., 2014b).
The influence of dust can also be felt indirectly by processing radiative energy via extinction. A particularly important example of this is photoelectric heating by dust grains; absorption of FUV photons by dust produces energetic free electrons that contribute to a heating of the gas component. In CHIMES, the volumetric heating rate, $\Gamma_{\rm PE}$, follows that of Wolfire et al. (2003) (their equations 19 and 20888This is an update of what is described in Richings et al. (2014a), described by Richings & Faucher-Giguère (2018).), and is given by
With the grain cross-section per unit volume the relevant quantity for all of these processes, we modulate them using the masses and sizes of grains traced through our live dust model. We describe their coupling in §3.6. A final dust-sensitive aspect to consider is the role of depletion. The same chemical elements that constitute dust also play an important cooling role in the ISM when existing in diffuse form, particularly through metal-line cooling. These processes are handled via per-ion cooling rates tabulated for COLIBRE and based upon Oppenheimer & Schaye (2013) (P25 equation 19101010Updated from Ploeckinger & Schaye (2020)). These rates are based on diffuse-phase metals, so it is necessary to assume a per-element depletion alongside the $\mathcal{DTG}_{\rm hyb}$, in lieu of a live dust model. These assume depletion in the fixed relative proportions of the Milky Way as inferred by Jenkins (2009). We discuss how we then adjust these rates for the case of live dust in §3.6, consistent with the depletions we calculate for our dust model (§3.8).
3.6 Coupling to COLIBRE hybrid cooling
@@
The evolutionary life-cycle model of dust described in §2 emphasises the influence of stars and gas in the ISM. Given this model, we can then in turn account for the influence of local dust properties on various radiative processes, such that dust may itself influence gas and star formation. In §3.5 we detail the relevant processes influenced by live dust grains. The relevant quantity for cooling processes linked to grains is the local dust cross-section per unit volume, $n_{\rm d}\sigma_{\rm d}$. The hybrid cooling module assumes a constant dust mix internally, through an ‘effective size’ translating to an $A_{V}/N_{\rm H}$ value (see §3.5). Having fixed attenuation curves parametrised solely by their normalisation in the $V$-band, $A_{V}$, means it is less complex and computationally expensive to compute heating and cooling rates. With mass-weighted average grain sizes of the MW ($a\sim 0.1$ µm, see e.g. Draine 2003) consistent with the effective size assumed for our OneSize model (Table 2, taken as our large-grain size in Fiducial and elsewhere) it is straightforward to simply scale extinctions and collisional cooling processes particle-by-particle using our modelled dust-to-gas ratio $\mathcal{DTG}_{\rm model}$. This is done via a rate scaling factor $s_{\rm dust}$ where $s_{\rm dust}=\mathcal{DTG}_{\rm model}/\mathcal{DTG}_{\rm MW}$, where we take $\mathcal{DTG}_{\rm MW}=6.6\times 10^{-3}$ (following the ISM grain model in Cloudy, Chatzikos et al. 2023).
@@
We note that for collisional processes (e.g. H2 formation) the $s_{\rm dust}$ scaling simply reflects the increased surface area per unit mass contributed by small grains. For radiative processes, however, this scaling may strictly depend upon an additional extinction efficiency, $\langle Q_{\rm ext}\rangle$, which can vary with wavelength and differs depending on the material and size properties of grains. While $V$-band attenuation $A_{V}$ is used as a standard normalisation for extinction curves, the processes relevant to the dust coupling described here (e.g. self shielding, photoelectric heating) actually depend on the FUV, ionising or Lyman-Werner photons. While extinction efficiencies of large grains, $\langle Q_{{\rm ext},L}\rangle$, dominate at optical wavelengths (e.g. Draine & Li, 2001), in the UV the small-grain extinction efficiency, $\langle Q_{{\rm ext},S}\rangle$, becomes comparable. For simplicity, we neglect the influence of $\langle Q_{\rm ext}\rangle$, assuming $\langle Q_{{\rm ext},S}\rangle/\langle Q_{{\rm ext},L}\rangle\approx 1$ and marginal differences between material species. We look into detailed extinction and attenuation properties of grain in more detail in upcoming studies post-processing COLIBRE galaxies using radiative transfer, and justify our assumptions around extinction efficiency further in Appendix A.
@@
The live dust model may also impact radiative cooling and heating processes more indirectly, through the different depletion of gas-phase elements, affecting gas-phase cooling channels (such as metal-line cooling and heating). The element-by-element cooling channels for diffuse metals in the hybrid cooling scheme accounts for the assumed P25 dust depletion, via a MW-like relative depletion pattern (Jenkins, 2009) that scales with $Z_{\rm gas}$ (see §3.8 for a comparison of equilibrium and live dust depletions). The fractions of metal elements in the diffuse- and dust-phase of enriched gas are handled implicitly, scaling the tabulated rates and leading to different element-by-element cooling and heating between the coupled and uncoupled cases for gas of a fixed thermodynamic state and metallicity, owing to the differences between our live dust model and the P25 dust.
@@
In Fig. 2 we show differences between dust, cooling and gas properties, comparing our live dust with the instantaneous P25 dust, both within the Fiducial run, and when incorporating the coupled physics effects when comparing to FidUncoupled runs. In the top row we see the influence of the dust model alone, by comparing the P25 dust to the live dust model within Fiducial. Here, we see that for relatively cool ($\log_{10}T/{\rm K}\lesssim 3.5$) and moderately dense ($\log_{10}n_{\rm H}/{\rm cm^{-3}}>-2$) gas, the P25 model tends to yield lower $\mathcal{DTG}$ ratios (panel a), though with small, exceptional regions (around $\log_{10}n_{\rm H}/{\rm cm^{-3}}\approx-0.5$, $\log_{10}T/{\rm K}\approx 2.5$ and $\log_{10}n_{\rm H}/{\rm cm^{-3}}\gtrapprox 2$). For hotter or more diffuse gas, the live dust model dominates, particularly due to the propagation of dust outside of the ISM and into the CGM and other phases that is assumed to be completely destroyed in the P25 model. This residual dust that avoids destruction by sputtering and direct annihilation in shocks is seen within Hii regions (horizontal sequence at higher density and $\log_{10}T/{\rm K}\approx 4$) and in the diffuse hot gas, where despite the high temperatures, dust destruction timescales can remain long enough due to the lower densities. We highlight the few cells in hot, denser gas showing evidence of live dust (around $\log_{10}T/{\rm K}\approx 7$, $\log_{10}n_{\rm H}/{\rm cm^{-3}}\approx 0$). These few particles are those caught in the simulation output between the feedback and dust evolution steps (where this dust is rapidly destroyed), due to the order of operations in COLIBRE. The total dust content of the Fiducial volume when using P25 or live dust is very similar (within 0.01 dex or $\approx 2\%$), though this level of consistency is coincidental, varying more in other volumes and environments. Recomputing the instantaneous P25 elemental cooling and heating rates of Fiducial using P25 dust (panels b and c respectively) yields complex differences, particularly higher metal-line cooling and heating rates in Hii regions (due to more diffuse metals) and heating in the densest ISM. However, in overall cosmic terms, the difference in cooling and heating rates per unit density ($\Lambda_{\rm cool}/n_{\rm H}$, $\Lambda_{\rm heat}/n_{\rm H}$) are marginal ($\lesssim 0.2\%$).
@@
The bottom row of Fig. 2 compares the live dust of Fiducial to the P25 dust of the FidUncoupled run, including the physical differences and run-run variations. To illustrate this, we first show the different gas distributions (panel d). Along with the variations in hot, diffuse gas owing to differing and stochastic feedback histories of each run, there are also differences in the ISM; most notably the existence of an excess of colder gas in the FidUncoupled run at $\log_{10}T/{\rm K}\lessapprox 100$. Taking this into account, we see that the local grain cross sections are generally higher in the live dust model, and by 0.53 dex (3$\approx$ 3$\times$ 3 ×) overall, excepting this cold sequence and a region of the WNM, at $\log_{10}n_{\rm H}/{\rm cm^{-3}}\approx-1$, $\log_{10}T/{\rm K}\approx 3.25$. This is attributable to our variable size distributions and abundance of small grains, particularly outside the coldest, densest gas. Another significant difference to highlight is that H2 extends to slightly warmer temperatures and lower densities in the coupled runs relative to the uncoupled ones (panel f). We attribute this to lower cooling rates in the dense ISM, paired with more small grains at intermediate density, boosting the local grain cross-section for H2 formation. Ultimately, the distribution of H2 in $\rho$-$T$ space leads to higher cosmic H2 mass densities (by 0.17 dex), boosting galactic H2 mass fractions. Altogether, this emphasises the complex differences that emerge from using a live dust model [^11] Alongside any stochastic run-to-run variation (Borrow et al., 2023)., where dust evolution is influencing gas thermodynamics throughout cosmic time.

### v2

@@ 3.5 Grain size evolution
Other processes have been identified that conserve the mass in dust, but modify the grain size distribution. Despite not affecting the grain mass directly, this has a knock-on effect on the evolution of grains, with the accretion and sputtering timescales being sensitive to grain size, with small grains exhibiting shorter timescales due to their higher surface to volume ratios.
@@ 3.5.1 Shattering
where the timescale normalisation value assumes a material density of 3 $\rm g\;cm^{-3}$} (intermediate between typical silicate and carbon grains), and $a_{L}$ is the radius of large grains (0.1 µm in our model). The Granato et al. (2021) prescription differs from previous applications in that the shattering does not shut off abruptly above a gas density of $n_{\rm H}=1\;{\rm cm^{-3}}$, but instead increases more gradually above this density. Note, as in sputtering, we are using the unmodulated density, $n_{\rm H}$ (as opposed to $n_{\rm H}^{\prime}$), as sputtering is a diffuse-gas process.
@@ 3.5.2 Coagulation
where the timescale normalisation, $\tau_{\rm co,\,0}$ is taken to be $27.1~{\rm Myr}$ and incorporates an assumed material density of 3 $\rm g\;cm^{-3}$} for grains, and $a_{S}$ is the radius of small grains (0.01 µm). We also assume a fixed coagulation velocity of $v_{\rm co}=0.2\;{\rm km\;s^{-1}}$. In the original prescription of Aoyama et al. (2017) this process is assumed to take place in molecular clouds, with a fixed density of $10^{3}~{\rm cm^{-3}}$. We adapt the Aoyama et al. (2017) equation to depend on an effective density, $n^{\prime}_{\rm H}$, i.e. the local simulated gas density boosted by a clumping factor to represent unresolved structure (see § 3.8), as with the other collisional processes assumed here. Equation 10 makes use of the higher densities afforded by the higher resolution and cold gas physics of our simulations, and does not depend on a sub-resolution criterion for molecular gas content. This allows us to maintain a self-consistent approach across the dust evolution processes. This is particularly important when dust is coupled to cooling and molecule formation processes (§3.7).
Despite our higher resolution and cold gas physics, we cannot resolve the interiors of molecular clouds, as addressed in section 3.3. The coagulation rates are therefore also subject to a variable clumping factor, explained further in §3.8. In addition to the clumping factor, we note that the coagulation timescale $\tau_{\rm co,\,0}$ is relatively uncertain, so calibration of $\tau_{\rm co,\,0}$, to match second-order dust properties, such as the size distribution of grains, is a reasonable approach. In particular, a lower $\tau_{\rm co,\,0}$ could be argued for, particularly due to the assumption of spherical grains; more prolate/oblate grains could increase grain-grain collision rates, and increase the growth of grains at fixed density. We will consider the option of boosting this coagulation rate in future work.
@@ 3.6 Grain transport & diffusion
3.7 Coupling to COLIBRE hybrid cooling and heating
@@
The evolutionary life-cycle model of dust described above emphasises the influence of stars and gas in the ISM. Given this model, we can in turn account for the influence of local dust properties on various heating and cooling processes, such that dust may in turn influence gas physics generally and even star formation. In §2.4 we detail the relevant processes influenced by live dust grains. The relevant quantity for cooling processes linked to grains is the local dust cross-section per unit volume, $n_{\rm d}\sigma_{\rm d}$. The hybrid cooling module assumes a constant dust mix internally, through an ‘effective size’ translating to an $A_{V}/N_{\rm H}$ value (see §2.4). Having fixed attenuation curves parametrised solely by their normalisation in the $V$-band, $A_{V}$, means it is less complex and computationally expensive to compute heating and cooling rates. With mass-weighted average grain sizes of the MW ($a\sim 0.1$ µm, see e.g. Draine 2003) consistent with the effective size assumed for our OneSize model (Table 1, taken as our large-grain size in Fiducial and elsewhere) it is straightforward to simply scale extinctions and collisional cooling processes particle-by-particle using our modelled dust-to-gas ratio $\mathcal{DTG}_{\rm model}$. This is done via a rate scaling factor $s_{\rm dust}$ where $s_{\rm dust}=\mathcal{DTG}_{\rm model}/\mathcal{DTG}_{\rm MW}$, where we take $\mathcal{DTG}_{\rm MW}=6.6\times 10^{-3}$ (following the ISM grain model in Cloudy, Chatzikos et al. 2023).
@@
We note that for collisional processes (e.g. H2 formation) the $s_{\rm dust}$ scaling simply reflects the increased surface area per unit mass contributed by small grains. For heating and cooling processes, however, this scaling may strictly depend upon an additional extinction efficiency, $\langle Q_{\rm ext}\rangle$, i.e. the ratio of the extincting cross section (the sum of scattering and absorption cross sections) to the physical cross section of grains (e.g. Weingartner & Draine, 2001). This can vary with wavelength and differs depending on the material and size properties of grains. While $V$-band attenuation $A_{V}$ is used as a standard normalisation for extinction curves, the processes relevant to the dust coupling described here (e.g. self shielding, photoelectric heating) actually depend on the FUV, ionising or Lyman-Werner photons. While extinction efficiencies of large grains, $\langle Q_{{\rm ext},L}\rangle$, dominate at optical wavelengths (e.g. Draine & Li, 2001), in the UV the small-grain extinction efficiency, $\langle Q_{{\rm ext},S}\rangle$, becomes comparable. For simplicity, we neglect the influence of $\langle Q_{\rm ext}\rangle$, assuming $\langle Q_{{\rm ext},S}\rangle/\langle Q_{{\rm ext},L}\rangle\approx 1$ and marginal differences between material species. We look into detailed extinction and attenuation properties of grain in more detail in upcoming studies post-processing COLIBRE galaxies using radiative transfer, and justify our assumptions around extinction efficiency further in Appendix B.
@@
The live dust model may also impact radiative cooling and heating processes more indirectly, through the different depletion of gas-phase elements, affecting gas-phase cooling channels (such as metal-line cooling and heating). The element-by-element cooling channels for gas-phase metals in the hybrid cooling scheme accounts for the assumed P25 dust depletion, via a MW-like relative depletion pattern (Jenkins, 2009) that scales with $Z_{\rm gas}$ (see §3.9 for a comparison of equilibrium and live dust depletions). The fractions of metal elements in the gas- and dust-phase of enriched gas are handled implicitly, scaling the tabulated rates and leading to different element-by-element cooling and heating between the coupled and uncoupled cases for gas of a fixed thermodynamic state and metallicity, owing to the differences between our live dust model and the P25 dust.
In Fig. 2 we show differences between dust, cooling and gas properties, comparing our live dust with the instantaneous P25 dust, both within the Fiducial run, and when incorporating the coupled physics effects when comparing to FidUncoupled runs. In the top row we see the influence of the dust properties input into the heating and cooling modules alone, by comparing the P25 dust to the live dust model within Fiducial. This shows how, for like physical conditions, the influence of dust on cooling processes differs across $n_{\rm H}$-$T$ space when live dust is included. The shading in this figure shows absolute pixel differences normalised by the maximum positive and negative difference pixel in each panel. This nuanced representation allows us to show where bulk properties differ in phase space. The overall distribution of dust depletion fractions in phase space can be seen more directly in upcoming § 4.2 with overall phase-spce distributions for COLIBRE presented in Schaye et al. (2025).
@@
In Fig. 2, we see that for relatively cool ($\log_{10}T/{\rm K}\lesssim 3.5$) and moderately dense ($\log_{10}n_{\rm H}/{\rm cm^{-3}}>-2$) gas, the P25 model tends to yield lower $\mathcal{DTG}$ ratios (panel a), though with small, exceptional regions (around $\log_{10}n_{\rm H}/{\rm cm^{-3}}\approx-0.5$, $\log_{10}T/{\rm K}\approx 2.5$ and $\log_{10}n_{\rm H}/{\rm cm^{-3}}\gtrapprox 2$). For hotter or more diffuse gas, the live dust model dominates, particularly due to the propagation of dust outside of the ISM and into the CGM and other phases that is assumed to be completely destroyed in the P25 model. This residual dust that avoids destruction by sputtering and direct annihilation in shocks is seen within Hii regions (horizontal sequence at higher density and $\log_{10}T/{\rm K}\approx 4$) and in the diffuse hot gas, where despite the high temperatures, dust destruction timescales can remain long enough due to the lower densities. We highlight the few cells in hot, denser gas showing evidence of live dust (around $\log_{10}T/{\rm K}\approx 7$, $\log_{10}n_{\rm H}/{\rm cm^{-3}}\approx 0$). These few particles are those caught in the simulation output between the feedback and dust evolution steps (where this dust is rapidly destroyed), due to the order of operations in COLIBRE. The total dust content of the Fiducial volume when using P25 or live dust is very similar (within 0.01 dex or $\approx 2\%$), though this level of consistency is coincidental, varying more in other volumes and environments. Recomputing the instantaneous P25 elemental cooling and heating rates of Fiducial using P25 dust (panels b and c respectively) yields complex differences, particularly higher metal-line cooling and heating rates in Hii regions (due to more gas-phase metals) and heating in the densest ISM. However, in overall cosmic terms, the difference in cooling and heating rates per unit density ($\Lambda_{\rm cool}/n_{\rm H}$, $\Lambda_{\rm heat}/n_{\rm H}$) are marginal ($\lesssim 0.2\%$).
@@
The bottom row of Fig. 2 compares the live dust of Fiducial to the P25 dust of the FidUncoupled run, including the physical differences and run-run variations. To illustrate this, we first show the different gas distributions (panel d). Along with the variations in hot, diffuse gas owing to differing and stochastic feedback histories of each run, there are also differences in the ISM; most notably the existence of an excess of colder gas in the FidUncoupled run at $\log_{10}T/{\rm K}\lessapprox 100$. Taking this into account, we see that the local grain cross sections are generally higher in the live dust model, and by 0.53 dex ($\approx 3\times$) overall, excepting this cold sequence and a region of the WNM, at $\log_{10}n_{\rm H}/{\rm cm^{-3}}\approx-1$, $\log_{10}T/{\rm K}\approx 3.25$. This is attributable to our variable size distributions and abundance of small grains, particularly outside the coldest, densest gas. Another significant difference to highlight is that H2 extends to slightly warmer temperatures and lower densities in the coupled runs relative to the uncoupled ones (panel f). We attribute this to lower cooling rates in the dense ISM, paired with more small grains at intermediate density, boosting the local grain cross-section for H2 formation. Ultimately, the distribution of H2 in $\rho$-$T$ space leads to higher cosmic H2 mass densities (by 0.17 dex), boosting galactic H2 mass fractions. Altogether, this emphasises the complex differences that emerge from using a live dust model [^13] Alongside any stochastic run-to-run variation (Borrow et al., 2023)., where dust evolution is influencing gas thermodynamics throughout cosmic time.